(GDF Deflator) The price of mumble rap records was $2 in 2015 and $2.50 in 2018. The number of these records produced was 150 in 2015 and rose by 30% in 2018. Beige couches were $200 in 2018, and $210 in 2015. There were 20 manufactured in 2015, but only 18 in 2018.

a) Calculate real GDP growth in percentages and dollars using both base years.

b) Calculate the GDP deflator using both base years

I'm struggling with these questions for an hour now. If someone could help me answer these questions, that would be nice. Thank you

to use Ginger
Limited mode
these questions for an hour
×
Answer #1

Given,

Price of mumble rap in 2015 = $2

Price of mumble rap in 2018 = $2.50

Quantity of mumble rap in 2015 = $150

Quantity of mumble rap in 2018 = rose by 30% = 150 + 30% of 150 = 150 + 45 = 195

Price of Beige couches in 2015 = $210

Price of Beige couches in 2018 = $200

Quantity of Beige couches in 2015 = $20

Quantity of Beige couches in 2018 = $18

a) real GDP growth using 2015 as base year = (real gdp in 2018 - real gdp in 2015 ) / real gdp in 2015

real GDP growth using 2018 as base year = (real gdp in 2018 - real gdp in 2015 ) / real gdp in 2018

Real GDP ( using 2015 as base year - means using price of year 2015)

Real GDP in 2015 = 2*150 + 210*20 = 4500

Real GDP in 2018 = 2*195 + 210*18 = 4170

real GDP growth using 2015 as base year = 4170 - 4500 / 4500 = - 7.33% (percentage)

real GDP growth using 2015 as base year = 4170 - 4500 = -330 ( in dollar terms)

Real GDP ( using 2018 as base year - means using price of year 2018)

Real GDP in 2015 = 2.50*150 + 200*20 = 4375

Real GDP in 2018 = 2.50*195 + 200*18 = 4087.5

real GDP growth using 2018 as base year = 4087.5 - 4375 / 4087.5 = - 7.03%

real GDP growth using 2018 as base year = 4087.5 - 4375 = -$287.5 ( in dollar)

b) GDP deflator = Nominal GDP / real GDP

GDP deflator in 2015 (using 2015 as base year) = 100 ( as both nominal and real gdp in 2015 is same when 2015 is base year)

GDP deflator in 2018 (using 2015 as base year) = nominal GDP in 2018 / real GDP in 2018

Nominal GDP in 2018 = 2.50*195 + 200*18 = 4087.5

real GDP in 2018 ( using 2015 as base year) =4170

GDP deflator in 2018 (using 2015 as base year) = (4087.5 / 4170)*100 = 98.02

GDP deflator (using 2018 as base year) = nominal GDP in 2018 / real GDP in 2018

GDP deflator in 2018 (using 2018 as base year) = 100

GDP deflator in 2015 (using 2018 as base year) = nominal GDP in 2015 / real gdp in 2015

Nominal GDP in 2015 = 4500

real GDP in 2015 ( using 2018 as base year) = 4375

GDP deflator in 2015 ( using 2018 as base year) = (4500 /4375)*100 = 102.86